OpenStudy (anonymous):

Please Help! A triangle has sides of and 3. Which could not be the length of the third side if it is a right triangle? Which of the following equations is of a parabola with a vertex at (0, -3)? Find the distance between the points (-5, 0) and (-4, 1). y(y + 4)(y - 6) = 0 is a quadratic equation. What is the x-coordinate of the vertex of the parabola whose equation is y = 3x2 + 12x + 5? What is the radius of the circle whose equation is x2 + y2 = 9?

OpenStudy (solomonzelman):

re-write your question.

OpenStudy (anonymous):

A triangle has sides of square route 2 and 3. Which could not be the length of the third side if it is a right triangle? i got square route of 11 but idk if its right @SolomonZelman

OpenStudy (solomonzelman):

\[\sqrt{2}^2+3^{3}=2+9=~~~~->~~~~\sqrt{11}\]OR\[3^3-\sqrt{2}^2=9-2=....~~~~~->~~~~\sqrt{7}\]

OpenStudy (anonymous):

ah, okay thank you sir! @SolomonZelman

OpenStudy (solomonzelman):

tnx for the sir :) yw !